A scientist wants to know how many fish are in a pond. He captures 15 and tags them and re-releases back into the pond. The next day he captures 15 more fish. Two of these have tags. How many fish are in the pond?
I have 4 answers to choose from, but do not know how to set up the problem in order to get the answer. Thanks

Answer by checkley77(12844)

15/x=2/15 cross multiply
2x=15*15
2x=225x=225/2
x=112.5 fish in the pond.